### v2

# Overcoming Common Flaws in the Evaluation of Selective Classification Systems

Abstract
Selective Classification, wherein models can reject low-confidence predictions, promises reliable translation of machine-learning based classification systems to real-world scenarios such as clinical diagnostics. While current evaluation of these systems typically assumes fixed working points based on pre-defined rejection thresholds, methodological progress requires benchmarking the general performance of systems akin to the $\mathrm{AUROC}$ in standard classification. In this work, we define 5 requirements for multi-threshold metrics in selective classification regarding task alignment, interpretability, and flexibility, and show how current approaches fail to meet them. We propose the Area under the Generalized Risk Coverage curve ($\mathrm{AUGRC}$), which meets all requirements and can be directly interpreted as the average risk of undetected failures. We empirically demonstrate the relevance of $\mathrm{AUGRC}$ on a comprehensive benchmark spanning 6 data sets and 13 confidence scoring functions. We find that the proposed metric substantially changes metric rankings on 5 out of the 6 data sets.
1 Introduction
Selective Classification (SC) is increasingly recognized as a crucial component for the reliable deployment of machine learning-based classification systems in real-world scenarios, such as clinical diagnostics [Dvijotham et al., 2023, Leibig et al., 2022, Dembrower et al., 2020, Yala et al., 2019]. The core idea of SC is to equip models with the ability to reject low-confidence predictions, thereby enhancing the overall reliability and safety of the system [Geifman and El-Yaniv, 2017, Chow, 1957, El-Yaniv et al., 2010]. This is typically achieved through three key components: a classifier that makes the predictions, a confidence scoring function (CSF) that assesses the reliability of these predictions, and a rejection threshold $\tau$ that determines when to reject a prediction based on its confidence score.
In evaluating SC systems, two primary concepts are essential: risk and coverage. Risk expresses the classifier’s error potential and is typically measured by its misclassification rate. Coverage, on the other hand, indicates the proportion of instances where the model makes a prediction rather than rejecting it. An effective SC system aims to minimize risk while maintaining high coverage, ensuring that the model provides accurate predictions for as many instances as possible.
Current evaluation of SC systems often focuses on fixed working points defined by pre-set rejection thresholds[Geifman and El-Yaniv, 2017, Liu et al., 2019, Geifman and El-Yaniv, 2019]. For instance, a common evaluation metric might be the selective risk at a given coverage of $70\%$ [Geifman and El-Yaniv, 2019], which communicates the potential risk associated with a specific confidence score $\mathrm{c}$ and selection threshold $\tau$ to a patient or end-user. While this method is useful for communicating risk in specific instances, it does not provide a comprehensive evaluation of the system’s overall performance. In standard classification, this is analogous to the need for the Area Under the Receiver Operating Characteristic (AUROC) curve rather than evaluating sensitivity at a specific specificity, which can be highly misleading when assessing general classifier capabilities [Maier-Hein et al., 2022]. The AUROC provides a holistic view of the classifier’s performance across all possible thresholds, thereby driving methodological progress. Similarly, SC requires a multi-threshold metric that aggregates performance across all rejection thresholds to fully benchmark the system’s capabilities.
Several current approaches attempt to address this need for multi-threshold metrics in SC. The Area Under the Risk Coverage curve ($\mathrm{AURC}$) is the most prevalent of these Geifman et al. [2018]. However, we demonstrate in this work that $\mathrm{AURC}$ has significant limitations as it fails to adequately translate the risk from specific working points into a meaningful aggregated evaluation score. This shortcoming hampers the ability to benchmark and improve SC methodologies effectively.
In our work, we aim to bridge this gap by providing a robust evaluation framework for SC. Our contributions are as follows:
Refined Task Formulation: We are the first to provide a comprehensive SC evaluation framework, explicitly deriving meaningful task formulations for different evaluation and application scenarios such as working point versus multi-threshold evaluation.
Formulation of Requirements: We define five critical requirements for multi-threshold metrics in SC, focusing on task suitability, interpretability, and flexibility. We assess current multi-threshold metrics against our six requirements and demonstrate their shortcomings.
Proposal of AUGRC: We introduce the Area Under the Generalized Risk Coverage ($\mathrm{AUGRC}$) curve, a new metric designed to overcome the flaws of current multi-threshold metrics for SC. $\mathrm{AUGRC}$ meets all five requirements, providing a comprehensive and interpretable measure of SC system performance.
Empirical Validation: We empirically demonstrate the relevance and effectiveness of $\mathrm{AUGRC}$ through a comprehensive benchmark spanning 6 datasets and 13 confidence scoring functions.
In summary, our work presents a significant advancement in the evaluation of SC systems, offering a more reliable and interpretable metric that can drive further methodological progress in the field.
Figure 1: The AUGRC metric based on Generalized Risk overcomes common flaws in current evaluation of Selective classification (SC). a) Refined task definition for SC. Analogously to standard classification, we distinguish between holistic evaluation for method development and benchmarking using multi-threshold metrics versus evaluation of specific application scenarios at pre-determined working points. The current most prevalent multi-threshold metric in SC, $\mathrm{AURC}$, is based on Selective Risk, a concept for working point evaluation that is not suitable for aggregation over rejection thresholds (red arrow). To fill this gap, we formulate the new concept of Generalized Risk and a corresponding metric, $\mathrm{AUGRC}$ (green arrow). b) We formalize our perspective on SC evaluation by identifying five key requirements for multi-threshold metrics and analyze how previous metrics fail to fulfill them. Abbreviations, CSF: Confidence Scoring Function, AU(G)RC: Area Under the (Generalized) Risk Coverage curve.
2 Refined Task Formulation
Selective Classification (SC) systems consist of a classifier $m:\mathcal{X}\rightarrow\mathcal{Y}$, which outputs a prediction and a confidence scoring function (CSF) $g:\mathcal{X}\rightarrow\mathbb{R}$, which outputs a confidence score associated with the prediction. Assuming a supervised training setup, let $\{(x_{i},y_{i})\}_{i=1}^{N}$ be a dataset containing $N$ independent samples from the source distribution $P(X,Y)$ over $\mathcal{X}\times\mathcal{Y}$. Given a rejection threshold $\tau$, the model prediction is accepted only if the corresponding score is larger than $\tau$:
$(m,g)(x)\coloneqq\begin{cases}m(x),&\text{if }g(x)\geq\tau\\ \text{reject},&\text{otherwise}\end{cases}$ (1)
Given an error function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$, the overall risk of the classifier $m$ is given by $R(m)\coloneqq\frac{1}{N}\sum_{i=1}^{N}\ell(m(x),y)$. The error function thereby contains the measure of classification performance suitable for the task at hand. Commonly, SC literature assumes a 0/1 error corresponding to the failure indicator variable $Y_{\text{f}}$ with $y_{\text{f},i}(x_{i},y_{i},m)\coloneqq\mathbb{I}[y_{i}\neq m(x_{i})]$. In this case, the overall risk represents the probability of misclassification $P(Y_{\text{f}}=1)$. By rejecting low-confidence predictions, the selective classifier can decrease the risk of misclassifications at the cost of letting the classifier make predictions on only a fraction of the dataset. This fraction is denoted as the $\text{coverage}\coloneqq P(g(x)\geq\tau)$, with the respective empirical estimator $\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(g(x_{i})\geq\tau)$. Evaluation of SC systems often focuses on preventing "silent", i.e. undetected failures for which both $y_{\text{f},i}=1$ and $g(x_{i})\geq\tau$.
Deploying a SC system in practice requires the selection of a fixed rejection threshold $\tau$. In the following refined evaluation protocol, we distinguish between the well-established application-specific task formulation where SC models are evaluated at individual working points (Section 2.1) and SC method development where evaluation is independent of individual working points (Section 2.2).
2.1 Evaluating SC systems in applied settings
Current Selective Classification evaluation commonly reports the selective risk of the system at a given coverage (“risk@coverage”), which implies a pre-determined cutoff on the risk or coverage (e.g. "maximum risk of 5%") leading to a working point of the system defined by a rejection threshold $\tau$. The selective risk is defined as:
$\mathrm{Selective\ Risk}_{(m,g)}(\tau)\coloneqq\frac{\sum_{i=1}^{N}\ell(m(x_{i}),y_{i})\cdot\mathbb{I}(g(x_{i})\geq\tau)}{\sum_{i=1}^{N}\mathbb{I}(g(x_{i})\geq\tau)}$ (2)
This formulation only considers accepted predictions with $c>\tau$. For a binary failure error, this risk is an empirical estimator for the conditional probability $P(Y_{\text{f}}=1|g(x)\geq\tau)$. Thus, this metric effectively communicates the risk of a "silent" failure for a specific prediction of classifier $m$ that has been accepted ($c>\tau$) at a pre-defined threshold $\tau$. This information may be useful in applied medical settings, e.g. to communicate the risk of misclassification for a patient p whose associated prediction has been accepted by a given SC system ($c_{p}>\tau$). Aside from the selective risk, other performance measurements for working point selection include for example Classification quality and Rejection quality [Condessa et al., 2017].
2.2 Evaluating SC systems for method development and benchmarking
Evaluating the general performance of a Selective Classification system requires a multi-threshold metric for a more comprehensive view compared to a working point analysis based on a fixed singular rejection threshold, analogous to binary classification tasks, where the Area Under the Receiver Operating Characteristic (AUROC) is used to assess the entire system performance across multiple classifier thresholds. However, the working point analysis can not be simply translated to multi-threshold aggregation, as it is based on Selective Risk, which only considers the risk w.r.t accepted predictions, assuming that a specific selection has already occurred ($P(Y_{\text{f}}=1\mid g(x)\geq\tau)$). This evaluation ignores rejected cases and thus contradicts the holistic assessment, where we are interested in the general risk for any prediction causing a silent failure independent of the future selection decision of individual predictions. This holistic assessment is reflected by the joint probability $P(Y_{\text{f}}=1,g(x)\geq\tau)$, which reflects the risk of silent failure for any prediction processed by the system including cases that are potentially rejected in the future.
To address the discrepancy between the need for holistic assessment versus the selective risk’s focus on already selected cases, we formulate the generalized risk:
$\mathrm{Generalized\ Risk}{(m,g)}(\tau)\coloneqq\frac{1}{N}\sum_{i=1}^{N}\ell(m(x_{i}),y_{i})\cdot\mathbb{I}[g(x_{i})\geq\tau].$ (3)
This metric reflects the joint probability of misclassification and acceptance by the confidence score threshold. By aggregating this risk over multiple rejection thresholds, we can evaluate the overall performance of the SC system.
2.3 Requirements for Selective Classification multi-threshold metrics
Based on the refined task definition given above, we can formulate concrete five requirements R1-R5 for multi-threshold metrics in Selective Classification:
R1: Task Alignment. The general goal in SC is to prevent silent failures either by preventing failures in the first place (via classifier performance), or by detecting the remaining failures (via CSF ranking quality). As argued in Jäger et al. [2023], it is crucial to jointly evaluate both aspects of an SC system, since the choice of CSF generally affects the underlying classifier. R2: Monotonicity. The metric should be monotonic w.r.t both evaluated factors stated in R1, i.e. improving on one of the two factors while keeping the other one fixed results in an improved metric value. Note that R2 does not make assumptions about how the two factors are combined, but it represents a minimum requirement for meaningful comparison and optimization of the metric. R3: Ranking Interpretability. Interpretability is a crucial component of a metric [Maier-Hein et al., 2022]. AUROC is the de facto standard ranking metric for binary classification tasks, providing an intuitive assessment of ranking quality which is proportional to the number of permutations needed for establishing an optimal ranking. This can also be interpreted as the "probability of a positive sample having a higher score than a negative one". Ideally, the SC metric should follow this intuitive assessment of rankings. R4: CSF Flexibility. The metric should be applicable to arbitrary choices of CSFs. This includes external CSFs, i.e. scores that are not based directly on the classifier output. R5: Error Flexibility. Current SC literature largely focuses on 0/1 error (i.e. $1-\text{accuracy}$) in their risk computation. However, in many real-world scenarios, accuracy is not an adequate classification metric, such as in the presence of class imbalance and for pixel-level classification. Thus, the SC metric should be flexible w.r.t the choice of error function.
2.4 Current multi-threshold metrics in SC do not fulfill requirements R1-R5
AURC: Geifman et al. [2018] derive the $\mathrm{AURC}$ as the Area under the Selective Risk-Coverage curve. This metric is the most prevalent multi-threshold metric in SC [Geifman et al., 2018, Jäger et al., 2023, Bungert et al., 2023, Cheng et al., 2023, Zhu et al., 2023a, Varshney et al., 2020, Naushad and Voiculescu, 2024, Van Landeghem et al., 2024, 2023, Zhu et al., 2022, Xin et al., 2021, Yoshikawa and Okazaki, 2023, Ding et al., 2020, Zhu et al., 2023b, Galil and El-Yaniv, 2021, Franc et al., 2023, Cen et al., 2023, Xia and Bouganis, 2022, Cattelan and Silva, 2023, Tran et al., 2022, Kim et al., 2023, Ashukha et al., 2020, Xia et al., 2024]. For the 0/1-error, $\mathrm{AURC}$ can be expressed through the following integral:
$\mathrm{AURC}=\int_{0}^{1}P(Y_{\text{f}}=1|g(x)\geq\tau)\,\mathrm{d}P(g(x)\geq\tau)$ (4)
This integral effectively aggregates the Selective Risk (Equation 2) over all fractions of accepted predictions (i.e. coverage). However, as discussed in Section 2.2, the Selective Risk is not suitable for aggregation over rejection thresholds to holistically assess a SC system. This inadequacy effectively leads to an excessive over-weighting of high-confidence failures in $\mathrm{AURC}$ and in return to breaching the requirements of monotonicity (R2) and ranking interpretability (R3). We provide empirical evidence for these shortcomings on toy data (Section 3) and real-world data (Section 4.2). Further, current SC literature usually employs the 0/1 error function, even though the related Accuracy metric is well known to be an unsuitable classification performance measure for many applications. For example, Balanced Accuracy is used to address class imbalance and metrics such as the Dice Score are employed for segmentation. Moving beyond the 0/1 error yields a higher variance in distinct error values and may thus amplify the shortcoming of over-weighting high-confidence failures. As more sophisticated error functions being are an important direction of future SC research, it is crucial to ensure the compatibility of metrics in SC evaluation.
e-AURC: Geifman et al. [2018] further introduce the excess-$\mathrm{AURC}$ as the difference to the $\mathrm{AURC}$ of an optimal confidence scoring function (denoted as $\mathrm{AURC}^{*}$)
$\mathrm{e\text{-}AURC}=\mathrm{AURC}-\mathrm{AURC}^{*}$ (5)
It relies on the intuition that subtracting the optimal $\mathrm{AURC}$ eliminates the contribution of the overall classification performance and collapses it to a pure ranking metric [Geifman et al., 2018, Galil et al., 2023, Jäger et al., 2023]. However, several works demonstrated that this intuition does not hold and that the $\mathrm{e\text{-}AURC}$ is still sensitive to the overall classification performance [Galil et al., 2023, Cattelan and Silva, 2023]. We attribute this deviation from the intended behavior to the shortcomings of the underlying $\mathrm{AURC}$, as we find the desired isolation of classifier performance in our improved metric formulation in Section 3. e-AURC further inherits the shortcomings of $\mathrm{AURC}$ w.r.t monotonicity (R2), ranking interpretability (R3), and error flexibility (R5).
[…]
Figure 2: The proposed AUGRC metric resolves shortcomings of AURC. All figures are based on rankings of predictions according to descending associated confidence scores induced by a CSF. All $\mathrm{AURC}$, $\mathrm{e\text{-}AURC}$, and $\mathrm{AUGRC}$ values are scaled by ×$\times 1000$. a) shows the contribution of an individual failure case on the $\mathrm{AURC}$ and $\mathrm{AUGRC}$ metrics depending on its ranking position (for technical details, see Section A.1.1). While $\mathrm{AUGRC}$ reflects the intuitive behavior of weighing the failure cases proportional to their ranking position, the $\mathrm{AURC}$ puts excessive weight on high-confidence failures. b-d) Toy example of three CSFs ranking 20 predictions to show how $\mathrm{AUGRC}$ resolves the broken monotonicity requirement (R2) of $\mathrm{AURC}$. Despite equal $\mathrm{AUROC_{f}}$ and equal $\mathrm{acc}$ in CSF-1 and CSF-2, the $\mathrm{AURC}$ improves. And $\mathrm{AURC}$ even improves in CSF-3, which features lower $\mathrm{AUROC_{f}}$ and lower $\mathrm{acc}$ compared to CSF-1. e-f) The corresponding risk-coverage curves reveal that the non-intuitive behavior of $\mathrm{AURC}$ is due to the excessive effect of the high-confidence failure of CSF-1 on the Selective Risk, which is resolved in the Generalized Risk.
3 Area under the Generalized Risk Coverage Curve
In Section 2.2, we illustrate that aggregating SC performance across working points requires to shift the perspective from the Selective Risk to the Generalized Risk (Equation 3) as an holistic assessment of the risk of silent failures for all predictions, before the rejection decision is made. We propose to evaluate SC methods via the Area under the Generalized Risk Coverage Curve ($\mathrm{AUGRC}$). For the binary failure error it becomes an empirical estimator of the following expression:
$\mathrm{AUGRC}=\int_{0}^{1}P(Y_{\text{f}}=1,\,g(x)\geq\tau)\,\mathrm{d}P(g(x)\geq\tau)$ (6)
This metric evaluates SC models in terms of the rate of silent failures averaged over working points and thus provides a practical measurement that is directly interpretable. It is bounded to the $[0,\nicefrac{{1}}{{2}}]$ interval, whereby lower $\mathrm{AUGRC}$ corresponds to better SC performance. The $\mathrm{AUGRC}$ is not subject to the shortcomings of the $\mathrm{AURC}$, and we can derive a direct relationship to $\mathrm{AUROC_{f}}$ and $\mathrm{acc}$ (the derivation and visualizations are shown in Equation 8 and Figure 5):
$\mathrm{AUGRC}=(1-\mathrm{AUROC_{f}})\cdot\mathrm{acc}\cdot(1-\mathrm{acc})+\frac{1}{2}(1-\mathrm{acc})^{2}$ (7)
To illustrate, when drawing two random samples, it corresponds to the probability that either both are failure cases or that one is a failure and has a higher confidence score than the non-failure. This is reflected by the second and first term of Equation 7, respectively. As an example, a CSF that outputs random scores yields $\mathrm{AUROC_{f}}=\frac{1}{2}$, and hence $\mathrm{AUGRC}=\frac{1}{2}(1-\mathrm{acc})$. The $\mathrm{AUGRC}$ for an optimal CSF ($\mathrm{AUROC_{f}}=1$) is given by the second term in Equation 7, hence subtracting the optimal SC performance yields a pure ranking measure re-scaled by the probability of finding a positive/negative pair. This overcomes the lack of interpretability of $\mathrm{AURC}$ and $\mathrm{e\text{-}AURC}$ (R3). Monotonicity (R2) is ensured since the partial gradients w.r.t. both $\mathrm{AUROC_{f}}$ and $\mathrm{acc}$ are always negative. Further, it can accommodate arbitrary classification metrics via the error function $\ell$ (R4) as well as arbitrary CSF (R5).
Figure 2a depicts the metric contribution of individual failure cases depending on their ranking position. This shows empirically how the excessive over-weighting of high-confidence failures in $\mathrm{AURC}$ is resolved by $\mathrm{AUGRC}$, and how the intuitive ranking assessment of $\mathrm{AUROC_{f}}$ is established (see Section A.1.1 for a detailed derivation). We further showcase how $\mathrm{AUGRC}$ resolves the broken monotonicity requirement (R2) of $\mathrm{AURC}$ in Figures 2 b-d. Despite equal $\mathrm{AUROC_{f}}$ and equal $\mathrm{acc}$ in CSF-1 and CSF-2, the $\mathrm{AURC}$ improves. And $\mathrm{AURC}$ even improves in CSF-3, which features lower $\mathrm{AUROC_{f}}$ and lower $\mathrm{acc}$ compared to CSF-1. Figures 2 e-f depict the associated risk-coverage curves and reveal that the non-intuitive behavior of $\mathrm{AURC}$ is due to the excessive effect of the high-confidence failure of CSF-1 on the Selective Risk, which is resolved in the Generalized Risk. In Section 4.2 we demonstrate implications of $\mathrm{AURC}$’s shortcomings on real-world data.
4 Empirical Study
We conduct an empirical study based on the existing FD-Shifts benchmarking framework [Jäger et al., 2023], which compares various CSFs across a broad range of datasets. The focus of our study is not on the performance of individual methods (CSFs) but rather on the ranking of methods based on $\mathrm{AUGRC}$ evaluation as compared to $\mathrm{AURC}$. We utilize the same experimental setup as in Jäger et al. [2023] with the addition of CSF scores based on temperature-scaled classifier logits. Extending the benchmark by recent SC methods, e.g. Feng et al. [2022], is an interesting direction of future work. A detailed overview of the datasets and methods used can be found in Appendix A.2. The code for reproducing our results and a PyTorch implementation of the $\mathrm{AUGRC}$ are available at: https://github.com/IML-DKFZ/fd-shifts.
4.1 Comparing Method Rankings of AUGRC and AURC
To study the relevance of $\mathrm{AUGRC}$, we investigate the changes induced by $\mathrm{AUGRC}$ on rankings of CSFs compared to $\mathrm{AURC}$ method rankings.
CSF Rankings with AUGRC substantially deviate from AURC rankings. Figure 3 illustrates the method ranking differences for all i.i.d. test datasets, showing changes in ranks across all CSFs. Notably, $\mathrm{AUGRC}$ induces changes in the top-3 methods (out of 13) on 5 out of 6 datasets. To ensure that these ranking changes are not due to variability in the test data, we evaluate the metrics on $500$ bootstrap samples from the test dataset and derive the compared method rankings based on the average rank across these samples ("rank-then-aggregate"). The size of each bootstrap sample corresponds to the size of the test dataset. Metric values for each bootstrap sample are averaged over 5 training initializations (2 for BREEDS and 10 for CAMELYON-17-Wilds). We analyze the robustness of the resulting method rankings for $\mathrm{AURC}$ and $\mathrm{AUGRC}$ separately, based on statistical significance. To that end, we perform pairwise CSF comparisons in form of a one-sided Wilcoxon signed-rank test [Wilcoxon, 1992] on the bootstrap samples. To control the family-wise error rate at a 5% significance level, we apply the Holm correction for multiple testing per metric and dataset, following [Wiesenfarth et al., 2021] (see Figure 7 for the results without correction for multiple testing). The resulting significance maps displayed in Figure 3 indicate stable method rankings for both $\mathrm{AURC}$ and $\mathrm{AUGRC}$. This suggests that the observed ranking differences are induced by the conceptual difference of $\mathrm{AUGRC}$ compared to $\mathrm{AURC}$ and, as a consequence, that the shortcomings of current metrics and respective solutions discussed in Section 2, are not only conceptually sound, but also highly relevant in practice. For the results in Figure 3, we select the DeepGamblers reward parameter and whether to use dropout for both metrics separately on the validation dataset (details in Table 1 and Table 2).
The shortcomings of AURC affect CSF comparison across datasets and distribution shifts. The method rankings for all datasets and distribution shifts, based on the original test datasets, are shown in Table 4, which also includes results for equal hyperparameters for $\mathrm{AURC}$ and $\mathrm{AUGRC}$. Results for the bootstrap-based method ranking analysis for distribution shifts are displayed in Figure 6. The substantial differences in method rankings across all datasets and distribution shifts underline the relevance of $\mathrm{AUGRC}$ for SC evaluation. The complete $\mathrm{AUGRC}$ results on the FD-Shifts benchmark are shown in Table 3.
In few cases, as shown in Figure 3, we observe that a CSF A is ranked worse than a neighboring CSF B even though CSF A’s ranks are significantly superior to those of CSF B based on the Wilcoxon statistic over the bootstrap samples. This discrepancy can occur if the rank variability in CSF A is larger than in CSF B. While this does not affect our conclusions regarding the relevance of the $\mathrm{AUGRC}$, it indicates that selecting a single best CSF for application based solely on method rankings should be approached with caution. We recommend closer inspection of the individual CSF performance in such cases, although this is not the focus of our study.
Figure 3: Substantial differences in method rankings for AUGRC and AURC. On 5 out of 6 datasets, the top-3 CSFs (out of 13 compared methods) change when employing the proposed $\mathrm{AUGRC}$ instead of $\mathrm{AURC}$. This demonstrates the practical relevance of the $\mathrm{AUGRC}$ metric for Selective Classification evaluation. CSFs are color-coded and sorted from top (best) to bottom (worst) by average rank based on $500$ bootstrap samples from the test dataset to ensure ranking stability. Ranking changes are reflected in changes in the color sequence and highlighted by red arrows. We assess the stability of the method rankings for each metric individually using one-sided Wilcoxon signed-rank tests based on the bootstrap samples at $5\%$ significance level with adjustment for multiple testing according to Holm. Adjacent to each ranking, we present the resulting significance maps for the pairwise CSF comparisons. These maps can be interpreted as follows: At each grid position $(x,y)$, filled entries indicate that metric values of CSF $y$ are ranked significantly better than those from CSF $x$ (across bootstrap samples), cross-marks indicate no significant superiority. An ideal ranking exhibits only filled entries above the diagonal.
4.2 Showcasing Implications of AURC Shortcomings on Real-world Data
Figure 4 gives an in-depth look at how the conceptual shortcomings of $\mathrm{AURC}$ affect method assessment on real-world data. The example uses the confidence like reservation score of the DeepGamblers method (DG-Res) [Liu et al., 2019] and the Monte Carlo Dropout-based predictive entropy (MCD-PE) [Gal and Ghahramani, 2016] as CSFs on the CIFAR-10 test dataset. Despite DG-Res having higher classification performance and ranking quality than MCD-PE, the $\mathrm{AURC}$ erroneously favors MCD-PE over DG-Res. This violates the monotonicity requirement (R2). This can be attributed to the excessive contributions of only few high-confidence failures, which aligns with the theoretical findings on failure contributions shown in Figure 2 (R3). In this example, the high-confidence failures are associated with high label ambiguity or incorrect labeling, suggesting that the $\mathrm{AURC}$ may exacerbate the influence of label noise in practice.
Figure 4: The conceptual shortcomings of AURC affects method assessment in practice. We illustrate the practical effects of excessive weight high-confidence failures in $\mathrm{AURC}$ by comparing the performance of two CSFs, DG-Res and MCD-PE, on the CIFAR-10 test dataset. (a) shows the coverage curves based on Selective Risk and Generalized Risk for both CSFs. The $\mathrm{AURC}$ violates the monotonicity requirement (R2) in practice, favoring MCD-PE despite a lower classification performance and ranking quality compared to DG-Res. (b) displays the images associated with the top-$k$ high-confidence failures. For DG-Res, the four failures correspond to the first four peaks in the Selective Risk curve, up to $\text{coverage}\approx 0.27$ (the total number of failures is 446). Only a few high-confidence failures significantly increase the $\mathrm{AURC}$. For both CSFs, the images associated with high-confidence failures exhibit high label ambiguity or are incorrectly labeled, indicating that the $\mathrm{AURC}$ may amplify the influence of label noise in practice. AURC and AUGRC values are scaled by ×$\times 1000$.
5 Conclusion
Despite the increasing relevance of SC for reliable translation of machine learning systems to real-world application, we find that the current metrics have limitations in providing the comprehensive assessment needed to guide the methodological progress of SC systems.
In this work, we establish a systematic SC evaluation framework, thereby promoting the adoption of more comprehensive, interpretable, and task-aligned metrics for comparative benchmarking of SC systems.
We find that none of the existing multi-threshold metrics, particularly the $\mathrm{AURC}$, meet the key requirements we identified for comprehensive SC evaluation, leading to deviations from intended and intuitive performance assessment behaviors. To address this, we introduce the $\mathrm{AUGRC}$ as a suitable metric for comprehensive SC method evaluation. Substantial differences in method rankings between $\mathrm{AURC}$ and $\mathrm{AUGRC}$, demonstrated through extensive empirical studies, highlight the importance of selecting the right SC metric. Thus, we propose the adoption of $\mathrm{AUGRC}$ for meaningful SC performance evaluation.
When evaluating specific applications for which certain risk or coverage intervals are known to be irrelevant, adaptations such as a partial AUGRC (analogous to the partial AUROC) may be considered. Further, investigating the relation between confidence ranking and calibration and on evaluating SC in settings where calibrated scores are of interest is an interesting direction of future work. Overall, our proposed evaluation framework provides a solid basis for future work in Selective Classification, including developing novel methods as well as analyzing the properties of individual methods.
Acknowledgments and Disclosure of Funding
This work was partly funded by Helmholtz Imaging (HI), a platform of the Helmholtz Incubator on Information and Data Science. We thank Lukas Klein, Maximilian Zenk and Fabian Isensee for insightful discussions and feedback.
Appendix A Appendix
A.1 Technical Details
A.1.1 AURC Failure Contribution
Let $\{\tau_{t}\}_{t=1}^{N}$ provide a total order on the predictions (relaxing the assumption to a partial order leads to a more complicated but very similar formulation). Then, for a rejection threshold $\tau_{t}$, a failure case at rank $t^{*}$ contributes to the Selective Risk (Equation 2) by $\frac{1}{t}$ if $t^{*}\leq t$. Averaging over all thresholds yields a contribution through the failure at rank $t^{*}$ to the $\mathrm{AURC}$ of $\frac{1}{N}\sum_{t=t^{*}}^{N}\frac{1}{t}$. For the Generalized Risk, the contribution at $t^{*}\leq t$ is always $\frac{1}{N}$, resulting in an overall contribution of $\frac{N-t^{*}-1}{N^{2}}$ to the $\mathrm{AUGRC}$. The failure contribution curves are displayed in Figure 2.
A.1.2 Relationship between AUGRC and failure AUROC
Let $Y_{\text{f}}=\mathbb{I}(m(x)\neq y)$ be a binary indicator variable for classification failures. Then, the Generalized Risk corresponds to the joint probability that a sample is accepted and wrongly classified. We can write the $\mathrm{AUGRC}$ as follows:
$\begin{split}\mathrm{AUGRC}&=\int P(Y_{\text{f}}=1,g(x)\geq\tau)\,\mathrm{d}P(g(x)\geq\tau)\\ &=\int(1-P(Y_{\text{f}}=0|g(x)\geq\tau))\cdot P(g(x)\geq\tau)\,\mathrm{d}P(g(x)\geq\tau)\\ &=\frac{1}{2}-\int P(g(x)\geq\tau|Y_{\text{f}}=0)P(Y_{\text{f}}=0)\,\mathrm{d}P(g(x)\geq\tau)\\ &\overset{(*)}{=}\frac{1}{2}-\text{acc}\Big{[}\frac{1}{2}\text{acc}+(1-\text{acc})\cdot\int P(g(x)\geq\tau|Y_{\text{f}}=0)\,\mathrm{d}P(g(x)\geq\tau|Y_{\text{f}}=1)\Big{]}\\ &=\frac{1}{2}-\text{acc}\Big{[}\frac{1}{2}\text{acc}+(1-\text{acc})\cdot\mathrm{AUROC_{f}}\Big{]}\\ &=(1-\mathrm{AUROC_{f}})\cdot\text{acc}(1-\text{acc})+\frac{1}{2}(1-\text{acc})^{2}\end{split}$ (8)
At (*), we use that $P(g(x)\geq\tau)=P(g(x)\geq\tau|Y_{\text{f}}=0)\cdot\text{acc}+P(g(x)\geq\tau|Y_{\text{f}}=1)\cdot(1-\text{acc})$. On choosing two random samples, the second term in the final equation represents the probability that both are failures and the first term represents the probability that one is a failure, the other is not, and that the failure has higher confidence score than the other.
The $\mathrm{AUGRC}$ is monotonic in $\mathrm{acc}$ and $\mathrm{AUROC_{f}}$ as both partial derivatives are negative $\forall\,\mathrm{acc}\in(0,1),\forall\,\mathrm{AUROC_{f}}\in[0,1]$:
$\frac{\partial\mathrm{AUGRC}}{\partial\mathrm{AUROC_{f}}}=-\mathrm{acc}\cdot(1-\mathrm{acc})$ (9)
$\frac{\partial\mathrm{AUGRC}}{\partial\mathrm{acc}}=2\cdot\mathrm{AUROC_{f}}\cdot\mathrm{acc}-\mathrm{acc}-\mathrm{AUROC_{f}}$ (10)
[…]
A.1.3 F1-AUC does not fulfill R2
For the optimal CSF, all failure cases are attributed to lower confidence scores than correct predictions. Hence, the precision $P(Y_{\text{f}}=0|g(x)\geq\tau)$ is one for coverages up to $P(Y_{\text{f}}=0)$ and decreases as $P(Y_{\text{f}}=0)/P(g(x)\geq\tau)$ above. Following the task formulation as defined in Malinin et al. [2021], we can derive the following analytical expression for the optimal F1-AUC:
$\begin{split}\text{F1-AUC}^{*}&=\int_{0}^{1}\text{F1}(\tau)\,\mathrm{d}P(g(x)\geq\tau)\\ &=\int_{0}^{P(Y_{\text{f}}=0)}\frac{2\cdot P(g(x)\geq\tau)}{P(Y_{\text{f}}=0)+P(g(x)\geq\tau)}\,\mathrm{d}P(g(x)\geq\tau)\\ &\quad+\int_{P(Y_{\text{f}}=0)}^{1}\frac{2\cdot P(Y_{\text{f}}=0)}{P(Y_{\text{f}}=0)+P(g(x)\geq\tau)}\,\mathrm{d}P(g(x)\geq\tau)\\ &=2\cdot\mathrm{acc}\cdot\Big{[}1+\ln\Big{(}\frac{1}{4}\cdot(1+\frac{1}{\mathrm{acc}})\Big{)}\Big{]}\end{split}$ (11)
F1-AUC increases with increasing accuracy up to $P(Y_{\text{f}}=0)\approx 0.56$, then it decreases, favoring models with lower classification performance. Thus, the F1-AUC does not fulfill the monotonicity requirement (R2).
A.2 Experiment Setup
A.2.1 Datasets and Methods
We compare the following CSFs: From the classifier’s logits we compute the Maximum Softmax Response (MSR), Maximum Logit Score (MLS), Predictive Entropy (PE), and the MLS based on temperature-scaled logits, for which a scalar temperature parameter is tuned based on the validation set [Guo et al., 2017]. Three predictive uncertainty measures are based on Monte Carlo Dropout [Gal and Ghahramani, ]: mean softmax (MCD-MSR), predictive entropy (MCD-PE), and mutual information (MCD-MI). We additionally include the DeepGamblers method [Liu et al., ], which learns a confidence like reservation score (DG-Res), ConfidNet Corbière et al. [2019], which is trained as an extension to the classifier, and the work of DeVries and Taylor [2018].
We evaluate SC methods on the FD-Shifts benchmark [Jäger et al., 2023], which considers a broad range of datasets and failure sources through various distribution shifts: SVHN [Netzer et al., 2011], CIFAR-10, and CIFAR-100 [Krizhevsky, ] are evaluated on semantic and non-semantic new-class shifts in a rotating fashion including Tiny ImageNet [Le and Yang, 2015]. Additionally, we consider sub-class shifts on iWildCam [Koh et al., 2021], BREEDS-Enity-13 [Santurkar et al., 2020], CAMELYON-17-Wilds [Koh et al., 2021], and on CIFAR-100 (based on super-classes) as well as corruption shifts based on Hendrycks and Dietterich [2019] on CIFAR-100. The data preparation and splitting are done as described in Jäger et al. [2023]. For the corruption shifts, we reduce the test split size to 75000 (subsampled within each corruption type and intensity level).
The following classifier architecture are used in the benchmark: small convolutional network for SVHN, VGG-13 [Simonyan and Zisserman, 2014] on CIFAR-10/100, ResNet-50 [He et al., 2016] on the other datasets.
If the distribution shift is not mentioned explicitly, we evaluate on the respective i.i.d. test datasets.
Our method ranking study focuses on the evaluation of CSF performance based on the existing FD-Shifts benchmark, hence we required no GPU’s for the analysis in Section. 4. As both $\mathrm{AURC}$ and $\mathrm{AUGRC}$ can be computed efficiently, (CPU) evaluation time for a single test set is less than a minute; evaluation on $500$ bootstrap samples on a single CPU core take around 3 hours.
A.2.2 Hyperparameters and Model Selection
The experiments are based on the same hyperparameters as reported in Table 4 in Jäger et al. [2023]. Based on the performance on the validation set, we choose the DeepGambler reward hyperparameter and whether to use dropout (for the non-MCD-based CSFs). For the former, we select from $[2.2,3,6,10]$ on Wilds-Camelyon-17, CIFAR-10, and SVHN, from $[2.2,3,6,10,15]$ on iWildCam and BREEDS-Entity-13, and from $[2.2,3,6,10,12,15,20]$ on CIFAR-100. When performing model selection based on the $\mathrm{AURC}$ metric, we obtain the same configurations as reported in Jäger et al. [2023]. When performing model selection based on the $\mathrm{AUGRC}$ metric, we obtain the parameters as reported in Tab. 1 and Tab. 2. For temperature scaling, we optimize the NLL on the validation set using the L-BFGS algorithm with $\mathrm{lr}=0.01$.
Table 1: Selected DeepGambler reward hyperparameter based on the $\mathrm{AUGRC}$ on the validation set for all confidence scoring functions trained with the DeepGamblers objective.
Method iWildCam BREEDS-Entity-13 Wilds-Camelyon-17 CIFAR-100 CIFAR-10 SVHN
DG-MCD-MSR 15 15 10 20 3 3
DG-PE 15 15 2.2 10 10 3
DG-Res 6 2.2 6 12 2.2 2.2
DG-TEMP-MLS 15 15 2.2 15 10 3
Table 2: Whether or not dropout-training has been selected based on the $\mathrm{AUGRC}$ on the validation set. This selection is only done for deterministic confidence scoring methods (no MCD). "1" denotes dropout training and "0" denotes training without dropout.
Method iWildCam BREEDS-Entity-13 Wilds-Camelyon-17 CIFAR-100 CIFAR-10 SVHN
MSR 0 0 1 0 1 1
MLS 0 0 1 1 1 1
PE 0 0 1 0 1 1
ConfidNet 1 0 1 1 1 1
DG-Res 0 1 0 0 0 1
Devries et al. 1 0 1 0 0 1
TEMP-MLS 0 0 1 0 1 1
DG-PE 1 0 0 0 0 1
DG-TEMP-MLS 1 0 0 0 0 1
A.3 AUGRC Computation Time
The main bottleneck of the $\mathrm{AUGRC}$ computation is the sort operation on the confidence scores. This is the same as for the $\mathrm{AURC}$ and there is no computational overhead compared to the $\mathrm{AURC}$. For a small benchmark of the computational time, we evaluate the $\mathrm{AURC}$, $\mathrm{AUGRC}$, and $\mathrm{AUROC_{f}}$ on 1000 random scores and failure labels. We obtain the following results (averaged over 5000 cases): $564\mu\!s\pm 2\mu\!s$ ($\mathrm{AURC}$), $562\mu\!s\pm 2\mu\!s$ ($\mathrm{AUGRC}$), $730\mu\!s\pm 13\mu\!s$ ($\mathrm{AUROC_{f}}$).
A.4 Additional Results
Table 3 shows the $\mathrm{AUGRC}$ results for the 13 compared CSFs across all datasets and distribution shifts. While the MSR baseline is not consistently outperformed across all experiments by any of the compared CSFs, MCD improves MSR scores in most settings. Temperature scaling does not exhibit consistent improvement of MLS scores.
Table 3: FD-Shifts benchmark results measured as AUGRC ×$\times 10^{3}$ (score range: [0, 500], lower is better $\downarrow$). The color heatmap is normalized per column, whereby whiter colors depict better scores. "cor" is the average over 5 intensity levels of image corruption shifts. AUGRC scores are averaged over 10 runs on CAMELYON-17-Wilds, over 2 runs on BREEDS, and over 5 runs on all other datasets. Abbreviations: ncs: new-class shift (s for semantic, ns for non-semantic), iid: independent and identically distributed, sub: sub-class shift, cor: image corruptions, c10/100: CIFAR-10/100, ti: TinyImagenet.
iWildCam BREEDS CAMELYON CIFAR-100 CIFAR-10 SVHN
study     iid sub    iid sub     iid sub     iid sub cor s-ncs ns-ncs     iid cor s-ncs ns-ncs     iid ns-ncs
ncs-data set                 c10 svhn ti    c100 svhn ti    c10 c100 ti
MSR     48.0 55.0    6.86 117     8.13 91.3     54.6 150 187 210 334 203    4.93 67.9 166 291 154    3.44 48.5 48.2 48.6
MLS     60.3 62.4    9.28 121     8.13 91.3     65.6 155 200 213 337 195    5.56 68.4 163 288 150    3.97 46.5 46.1 46.6
PE     48.2 55.2    6.87 116     8.13 91.3     56.4 150 188 210 333 202    4.91 67.2 165 290 152    3.44 47.9 47.5 48.0
MCD-MSR     42.8 54.3    6.89 116     5.64 95.1     53.1 146 176 210 336 202    4.56 60.4 166 294 156    3.40 47.3 47.4 47.9
MCD-PE     43.4 55.1    6.98 116     5.64 95.1     56.6 146 179 210 335 198    4.73 60.6 164 293 153    3.47 46.3 46.4 46.9
MCD-MI     43.7 58.1    7.28 118     6.13 101     59.4 148 182 211 333 197    4.99 64.9 167 302 161    3.55 47.0 47.7 48.2
ConfidNet     82.0 91.2    6.89 117     4.36 86.1     57.5 154 192 214 340 200    4.70 65.4 165 291 153    3.43 48.6 48.4 48.8
DG-MCD-MSR     43.9 59.6    6.31 112     3.47 149     52.5 145 175 210 336 203    4.75 60.8 168 296 159    3.38 47.6 47.7 48.0
DG-Res     66.5 63.8    9.39 122     3.46 124     64.3 230 194 215 330 197    4.40 63.8 167 287 151    4.09 46.8 46.2 46.6
Devries et al.     61.7 74.7    8.24 120     24.3 145     69.6 150 214 222 342 211    4.57 70.8 166 287 152    5.56 48.3 48.5 49.9
TEMP-MLS     48.2 55.1    6.86 116     8.13 91.3     54.8 150 187 210 333 203    4.92 67.6 166 291 153    3.44 48.3 47.9 48.4
DG-PE     52.0 69.7    6.33 114     3.46 123     55.2 151 186 210 329 197    4.09 62.9 167 288 154    3.41 48.9 48.4 48.7
DG-TEMP-MLS     52.2 69.9    6.35 114     3.46 123     55.0 222 184 211 331 199    4.09 63.1 167 288 155    3.41 49.0 48.5 48.8
Table 4: Comparing Rankings of AURC $\rightarrow$ $\alpha$ versus AUGRC $\rightarrow$ $\beta$. Differences in the method rankings between $\mathrm{AURC}$ and $\mathrm{AUGRC}$ demonstrate the relevance of the pitfalls of the $\mathrm{AURC}$ discussed in Section 2.4. Upper half: Model selection for dropout and DG hyperparameter was done for both metrics separately. Lower half: Model selection was done based on $\mathrm{AUGRC}$ for both $\alpha$ and $\beta$. The selected hyperparameters are reported in Appendix A.2.2. The color heatmap is normalized per column, whereby whiter colors depict better scores. "cor" is the average over 5 intensity levels of image corruption shifts. AUGRC scores are averaged over 10 runs on CAMELYON-17-Wilds, over 2 runs on BREEDS, and over 5 runs on all other datasets. Abbreviations: ncs: new-class shift (s for semantic, ns for non-semantic), iid: independent and identically distributed, sub: sub-class shift, cor: image corruptions, c10/100: CIFAR-10/100, ti: TinyImagenet.
iWildCam     BREEDS     CAMELYON     CIFAR-100     CIFAR-10     SVHN
study     iid     sub     iid     sub     iid     sub     iid     sub     cor     s-ncs     ns-ncs     iid     cor     s-ncs     ns-ncs     iid     ns-ncs
ncs-data set                                         c10     svhn     ti             c100     svhn     ti         c10     c100     ti
metric     $\alpha$ $\beta$    $\alpha$ $\beta$    $\alpha$ $\beta$    $\alpha$ $\beta$    $\alpha$ $\beta$    $\alpha$ $\beta$    $\alpha$ $\beta$    $\alpha$ $\beta$    $\alpha$ $\beta$    $\alpha$ $\beta$    $\alpha$ $\beta$    $\alpha$ $\beta$    $\alpha$ $\beta$    $\alpha$ $\beta$    $\alpha$ $\beta$    $\alpha$ $\beta$    $\alpha$ $\beta$    $\alpha$ $\beta$    $\alpha$ $\beta$    $\alpha$ $\beta$    $\alpha$ $\beta$
optimized both (w.r.t. dropout and DG reward parameter)
MSR     5 5    6 2    6 4    7 9    9 9    2 2    3 3    5 5    7 7    3 1    3 6    11 10    12 12    11 11    5 6    9 7    8 9    8 6    10 10    9 9    9 9
MLS     10 10    10 8    13 12    13 12    9 9    2 2    11 12    7 11    12 12    9 10    7 11    1 1    13 13    12 12    1 1    2 3    1 1    11 11    3 3    1 1    1 1
PE     6 6    7 5    8 6    6 4    9 9    2 2    10 7    9 5    11 9    10 1    12 4    6 8    10 10    9 9    4 4    6 6    2 3    8 6    7 7    6 6    6 6
MCD-MSR     1 1    1 1    3 7    3 4    6 6    6 7    2 2    2 2    2 2    1 1    9 8    9 8    3 4    2 2    5 6    12 12    10 12    4 2    5 5    5 5    5 5
MCD-PE     2 2    2 3    4 9    3 4    6 6    6 7    7 8    2 2    3 3    1 1    6 7    3 4    5 7    3 3    3 2    11 11    8 6    6 9    1 1    4 4    4 4
MCD-EE     3 3    3 6    5 10    3 4    8 8    6 6    8 9    2 2    4 4    6 8    7 8    6 6    7 9    1 1    2 2    9 7    2 3    8 10    2 2    3 3    3 3
ConfidNet     13 13    13 13    10 7    9 9    5 5    1 1    9 10    7 10    9 10    11 11    11 12    4 6    9 6    8 8    9 4    6 7    6 6    5 5    11 11    10 10    10 11
DG-MCD-MSR     4 4    9 7    1 1    1 1    4 4    12 13    1 1    1 1    1 1    3 1    10 8    11 10    6 8    4 4    10 13    13 13    13 13    1 1    6 6    7 7    7 6
DG-Res     12 12    11 9    12 13    12 13    1 1    9 11    12 11    13 13    10 11    12 12    1 2    2 2    8 3    7 7    13 10    3 1    5 2    12 12    4 4    2 2    2 1
Devries et al.     11 11    12 12    11 11    11 11    13 13    13 12    13 13    10 5    13 13    13 13    13 13    13 13    4 5    13 13    5 6    1 1    2 3    13 13    8 8    13 12    13 13
TEMP-MLS     7 6    8 3    8 4    7 4    9 9    2 2    5 4    5 5    8 7    3 1    3 4    10 10    11 11    10 10    5 6    8 7    6 6    7 6    8 8    8 8    8 8
DG-PE     8 8    4 10    7 2    10 2    1 1    10 9    6 6    11 9    6 6    6 1    2 1    5 2    1 1    5 5    11 10    4 3    10 9    2 3    12 12    10 10    10 10
DG-TEMP-MLS     8 9    4 11    2 3    2 2    1 1    10 9    4 5    12 12    5 5    8 8    3 3    6 5    1 1    6 6    11 10    5 3    12 11    3 3    13 13    12 12    12 11
optimized AUGRC (w.r.t. dropout and DG reward parameter)
MSR     5 5    4 2    7 4    8 9    9 9    2 2    3 3    7 5    7 7    4 1    5 6    11 10    12 12    11 11    5 6    9 7    8 9    8 6    10 10    9 9    9 9
MLS     10 10    8 8    13 12    13 12    9 9    2 2    11 12    10 11    12 12    10 10    8 11    1 1    13 13    12 12    1 1    2 3    1 1    11 11    3 3    1 1    1 1
PE     6 6    5 5    8 6    7 4    9 9    2 2    9 7    6 5    9 9    1 1    4 4    8 8    10 10    9 9    4 4    6 6    2 3    8 6    7 7    6 6    6 6
MCD-MSR     1 1    1 1    4 7    4 4    6 6    6 7    2 2    2 2    2 2    1 1    10 8    8 8    3 4    2 2    5 6    12 12    10 12    4 2    5 5    5 5    5 5
MCD-PE     2 2    2 3    5 9    4 4    6 6    6 7    7 8    2 2    3 3    1 1    7 7    4 4    5 7    3 3    3 2    11 11    8 6    6 9    1 1    4 4    4 4
MCD-EE     3 3    3 6    6 10    4 4    8 8    6 6    8 9    2 2    4 4    7 8    8 8    7 6    7 9    1 1    2 2    9 7    2 3    8 10    2 2    3 3    3 3
ConfidNet     13 13    13 13    10 7    10 9    5 5    1 1    10 10    10 10    10 10    11 11    12 12    5 6    9 6    8 8    9 4    6 7    6 6    5 5    11 11    10 10    10 11
DG-MCD-MSR     4 4    7 7    1 1    1 1    4 4    12 13    1 1    1 1    1 1    4 1    11 8    11 10    6 8    4 4    10 13    13 13    13 13    1 1    6 6    7 7    7 6
DG-Res     12 12    9 9    12 13    12 13    1 1    9 11    12 11    13 13    11 11    12 12    1 2    2 2    8 3    7 7    13 10    3 1    5 2    12 12    4 4    2 2    2 1
Devries et al.     11 11    12 12    11 11    11 11    13 13    13 12    13 13    5 5    13 13    13 13    13 13    13 13    4 5    13 13    5 6    1 1    2 3    13 13    8 8    13 12    13 13
TEMP-MLS     7 6    6 3    8 4    8 4    9 9    2 2    4 4    7 5    8 7    4 1    5 4    10 10    11 11    10 10    5 6    8 7    6 6    7 6    8 8    8 8    8 8
DG-PE     8 8    10 10    1 2    2 2    1 1    10 9    5 6    9 9    6 6    7 1    1 1    3 2    1 1    5 5    11 10    4 3    10 9    2 3    12 12    10 10    10 10
DG-TEMP-MLS     9 9    11 11    3 3    2 2    1 1    10 9    6 5    12 12    5 5    9 8    3 3    5 5    1 1    6 6    11 10    5 3    12 11    3 3    13 13    12 12    12 11
Figure 6: Method ranking analysis for distribution shifts. This figure displays the results for the method ranking analysis for $\mathrm{AURC}$ and $\mathrm{AUGRC}$ for evaluation under distribution shifts, analogous to Figure 3. The subfigures are titled as follows: "training dataset" $\rightarrow$ "evaluation dataset". Details on the distribution shifts are given in Section A.2.1. CSFs are color-coded and sorted from top (best) to bottom (worst) by average rank based on $500$ bootstrap samples from the test dataset to ensure ranking stability. Ranking changes are reflected in changes in the color sequence and highlighted by red arrows. We assess the stability of the method rankings for each metric individually using one-sided Wilcoxon signed-rank tests based on the bootstrap samples at $5\%$ significance level with adjustment for multiple testing according to Holm. Adjacent to each ranking, we present the resulting significance maps for the pairwise CSF comparisons. These maps can be interpreted as follows: At each grid position $(x,y)$, filled entries indicate that metric values of CSF $y$ are ranked significantly better than those from CSF $x$ (across bootstrap samples), cross-marks indicate no significant superiority. Abbreviations: ncs: new-class shift (s for semantic, ns for non-semantic).
Figure 7: Method ranking analysis without correction for multiple testing. In contrast to the results shown in Figure 3, which include the Holm correction for multiple testing, this figure displays the results without correction. On 5 out of 6 datasets, the top-3 CSFs (out of 13 compared methods) change when employing the proposed $\mathrm{AUGRC}$ instead of $\mathrm{AURC}$. This demonstrates the practical relevance of the $\mathrm{AUGRC}$ metric for Selective Classification evaluation. CSFs are color-coded and sorted from top (best) to bottom (worst) by average rank based on $500$ bootstrap samples from the test dataset to ensure ranking stability. Ranking changes are reflected in changes in the color sequence and highlighted by red arrows. We assess the stability of the method rankings for each metric individually using one-sided Wilcoxon signed-rank tests, each with a $5\%$ significance level, based on the bootstrap samples. Adjacent to each ranking, we present the resulting significance maps for the pairwise CSF comparisons. These maps can be interpreted as follows: At each grid position $(x,y)$, filled entries indicate that metric values of CSF $y$ are ranked significantly better than those from CSF $x$ (across bootstrap samples), cross-marks indicate no significant superiority. An ideal ranking exhibits only filled entries above the diagonal.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The described pitfalls of existing metrics affect Selective Classification evaluation in general. Hence, we thoroughly discuss all common metrics in the field and demonstrate the relevance on a state-of-the-art benchmark.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The empirical study is based on an extensive, state-of-the-art, yet naturally limited benchmark. However, the scope is clearly defined in the paper.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: For the analytical relationships, we provide the assumptions and derivations.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The evaluation protocol as well as all experiment setups are described. Additionally, we provide the code for reproducing the results.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The code for reproducing the experimental results is published on github. All datasets used are freely accessible.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The details of the experiments setups are described in the paper.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We validate the significance of our results via statistical tests based on bootstrapping.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Details are mentioned in the Appendix.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: There are no conflicts with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We don’t see any potential negative societal impacts. We address enhanced reliability of real-world decision systems through the studied methods as a potential positive impact. However, we propose no new Selective Classification method in our work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our experiments do not involve any new datasets or classification models.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We added proper credit for all datasets and models used in our experiments.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide the code for our proposed evaluation protocol including an implementation of our proposed metric. Documentation is added for reproducibility.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer:[N/A]
  69. Justification: Not applicable.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Not applicable.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.